OpenStudy (anonymous):

WHAT WILL BE LENGTH OF LINE ab IN A TRIANGLE WHOSE ANGLE acb = 1 DEGREE, LINE ac= 1 KILOMETER, LINE, LINE bc = 1 KILOMETER

OpenStudy (campbell_st):

well it looks like the cosine rule, 2 sides and the included angle. I'd suggest you change the 1 kilometre measurements to metres to make the final answer easier to understand. here is a diagram  use \[AC^2 = 1000^2 + 1000^2 - 2 \times 1000 \times 1000 \times \cos(1)\]